OpenStudy (anonymous):

what is sin(t) if cos(t)=3/sqrt(11)?

OpenStudy (kropot72):

 Using the theorem of Pythagorus: \[(\sqrt{11})^{2}=3^{2}+x ^{2}\] \[x ^{2}=(\sqrt{11})^{2}-3^{3}=11-9=2\] \[x=\sqrt{2}\] Can you find sin(t) now?

OpenStudy (anonymous):

yes thank you!

OpenStudy (kropot72):

You're welcome :)